OpenStudy (nuccioreggie):

Jake throws a dart that hits the square shown below: A square is drawn. A circle of radius 8 units touches the sides of the square. What is the probability that the dart hits a point in the circle? 16% 21.5% 64% 78.5% I got D

OpenStudy (nuccioreggie):

@TheSmartOne please help me

OpenStudy (anonymous):

the idea is that it is the ratio of the areas

OpenStudy (nuccioreggie):

Ok is d correct?

OpenStudy (anonymous):

square has area \(16^2\) if i read it correctly

OpenStudy (anonymous):

circle has area \(64\pi\)

OpenStudy (anonymous):

calculate \[\frac{64\pi}{16^2}\] use a calculator

OpenStudy (nuccioreggie):

Ok here it is so u can get it better

OpenStudy (anonymous):

yeah i got D too

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=64pi%2F16^2

OpenStudy (anonymous):

go with D looks good

OpenStudy (nuccioreggie):

Ok can u help me with like 4 or 5 more please ill tag u ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

if i know them that is

OpenStudy (nuccioreggie):

Henry tossed a coin several times and recorded the number of heads and tails in the following table: Coin Experiment Side of Coin Number of Tosses Heads 12 Tails 18 What is the experimental probability of getting heads? 0.4 0.7 1.2 1.8

OpenStudy (anonymous):

experimental probability means "relative frequency: ratio of favorable to total

OpenStudy (anonymous):

total tosses?

OpenStudy (nuccioreggie):

I think c or b

OpenStudy (anonymous):

total tosses?

OpenStudy (anonymous):

yeah it has to be A or B must be less than one

OpenStudy (nuccioreggie):

Its A

OpenStudy (anonymous):

it is definitely \(\frac{12}{30}\) reduce i guess

OpenStudy (anonymous):

yeah \(0.4\)

OpenStudy (nuccioreggie):

A paper cup is dropped and its landing position is recorded. The cup can land on the side, on the open end, or on the closed end. The results of 20 trials are shown in the table below: A chart shows tally marks. Main title for the chart is Paper Cup Experiment. The title for the second column is Number of times occurred. The tally mark corresponding to open shows a set of 4 vertical lines cut by a slanting line. The tally mark corresponding to closed shows a set of 4 vertical lines cut by a slanting line. The tally mark corresponding to side shows 2 sets of 4 vertical lines cut by a slanting line. Based on the table, which of the following best compares the experimental probability of the cup landing on its open end with the experimental probability of the cup landing on its closed end? The probabilities are equal. The probability of landing on the open end is greater. The probability of landing on the closed end is greater. No conclusion can be made.

OpenStudy (nuccioreggie):

Its C i think

OpenStudy (anonymous):

i can't really understand that

OpenStudy (nuccioreggie):

Ok 1sec

OpenStudy (nuccioreggie):

Here it is

OpenStudy (anonymous):

open and closed are the same right? both five

OpenStudy (anonymous):

so equal

OpenStudy (nuccioreggie):

Yea i say. c

OpenStudy (anonymous):

not C, A

OpenStudy (nuccioreggie):

But its an extra one at the bottom

OpenStudy (anonymous):

let me look again

OpenStudy (anonymous):

no looks like 5 and 5 to me